# Leveraging Temporal Contextualization for Video Action Recognition

### v1

@@ Abstract
Pretrained vision-language models have shown effectiveness in video understanding. However, recent studies have not sufficiently leveraged essential temporal information from videos, simply averaging frame-wise representations or referencing consecutive frames. We introduce Temporally Contextualized CLIP (TC-CLIP), a pioneering framework for video understanding that effectively and efficiently leverages comprehensive video information. We propose Temporal Contextualization (TC), a novel layer-wise temporal information infusion mechanism for video that extracts core information from each frame, interconnects relevant information across the video to summarize into context tokens, and ultimately leverages the context tokens during the feature encoding process. Furthermore, our Video-conditional Prompting (VP) module manufactures context tokens to generate informative prompts in text modality. We conduct extensive experiments in zero-shot, few-shot, base-to-novel, and fully-supervised action recognition to validate the superiority of our TC-CLIP. Ablation studies for TC and VP guarantee our design choices. Code is available at https://github.com/naver-ai/tc-clip.
@@ 1 Introduction
Figure 1: Temporal information learning methods. Prior works considered temporal cues during the encoding process via (a) cross-frame attention [22, 32] with [CLS] token interactions or (b) temporal window expansion [33] by adding adjacent frame tokens to key-value pairs. However, the former lacks patch-level interactions, while the latter limits the range of temporal locality. (c) joint space-time attention allows full interactions across all tokens, but it is costly and incurs suboptimal performance (see Fig. 3). (d) Unlike prior approaches, our method aggregates pivotal tokens from a broader range yet efficiently for enhanced temporal integration into key-value pairs.
Pretrained large-scale Vision-Language Models (VLMs) have shown remarkable generalization capability in video understanding and have emerged as promising tools even for zero-shot or open-vocabulary recognition tasks [24, 10, 36]. However, pretraining task-specific models using video-text pairs pose significant challenges, primarily due to substantial computational costs as well as excessive expense for annotated video-text data [31, 34]. Consequently, recent studies in video understanding [30, 22, 13, 32, 26, 33, 5, 9] have shifted their focus towards employing image-based VLMs such as CLIP [24] via fine-tuning—aligning video representations with text embeddings derived from category names.
Despite the successful generalization of CLIP in video recognition, existing approaches fail to fully exploit temporal information in the video feature learning process. Specifically, the token interactions remain restrictive since they either rely on simple averaging of frame-wise representations [26], or attempt to gather information in the temporal axis only through class tokens [32, 22] as illustrated in Fig. 1(a). Although VCLIP [33] incorporates patch-level details by bringing keys and values from neighborhood frames in its self-attention operation, as in Fig. 1(b), its temporal scope is still limited to a local window. Such naïve computation for video representations makes models biased towards static information in their representation learning (e.g., objects and backgrounds) and hampers learning temporal information (e.g., motion and temporal variations). To ensure the global interactions of patch tokens in a spatio-temporal domain, one possible option is to consider every patch token from all frames as a reference during the encoding process as illustrated in Fig. 1(c).
Unfortunately, such a straightforward extension for temporally global interactions in VLMs pretrained with short image-text pairs witnesses extrapolation challenges [23, 4] due to the huge number of tokens in each video. We have observed that a naïve extension of sequence length along the temporal axis degrades its discriminability substantially, as shown in Fig. 2. The joint space-time attention model distributes its attention broadly across the patches but fails to concentrate on informative tokens to recognize actions, incurring suboptimal performance compared to the frame-wise attention baseline. Moreover, this approach suffers from heavy computational overhead due to numerous redundant and similar tokens often corresponding to background regions.
This paper presents Temporally Contextualized CLIP (TC-CLIP), a novel paradigm for extending CLIP to the video domain by encoding holistic video information through advanced temporal analysis. Specifically, our Temporal Contextualization (TC) pipeline summarizes global-range action cues into a small set of tokens for reference during the encoding process. Based on the simple idea by utilizing [CLS] query for self-attention, we could extract informative tokens from each frame and group them contextually. These context tokens act as additional key-value pairs for attention operations in the subsequent layer, presumably serving as a temporal bridge that conveys the video-level context throughout the encoding process. Our preliminary study, shown in Fig. 3, implies that existing methods shown in Fig. 1 offer minimal improvement over the frame-wise attention, highlighting the need for enhanced token interactions.
@@
Figure 2: Pitfall of joint space-time attention. Extending CLIP’s temporal sequence length degrades attention quality, presumably because it was not trained on such long sequences.
Figure 3: Quantitative analysis on temporal modelings. We compare the action recognition performance in the few-shot settings on diverse datasets. All existing methods fall behind our method.
@@
Furthermore, we propose the Video-conditional Prompting (VP) module for the generation of instance-level textual prompts based on produced context tokens. The VP module is composed of cross-attention operations that adopt learnable text prompts as queries and visual temporal tokens as keys and values for injecting video instance representations into video-conditional textual prompts. This strategy alleviates the lack of textual semantics in video action recognition datasets, where category names are the only descriptions of actions (e.g., skateboarding, skydiving, ski jumping), without detailed narratives about acting subjects, observed objects and environment, etc.
[…]
2 Related Work
CLIP for video action recognition. Recent video understanding studies [30, 13, 22, 32, 26, 33, 20, 5, 9] focus on adapting language-image pre-trained models [24, 10, 36]. As a pioneering work, ActionCLIP [30] introduces CLIP [24] into video recognition by adding temporal attention layers after the backbone in a late-fusion fashion. To consider global-range temporal cues during feature encoding, X-CLIP [22] and Vita-CLIP [32] perform cross-frame attention, promoting [CLS] token interaction before spatial attention at each layer. On the other hand, there are lines of work that fully fine-tune CLIP with minimal architectural modifications [26, 33, 5, 9]. ViFi-CLIP [26], a simple fine-tuning method with CLIP without additional temporal modules, generalizes well in open-vocabulary learning. Open-VCLIP [33] expands the temporal view of self-attention by adding keys and values of neighboring frames. Despite the generalization capacity of recent models, we argue they did not adequately leverage temporal information during the video feature encoding process, as they either rely on averaging frame-wise representations [26], [CLS] token interactions [22, 32], or considering only consecutive frames [33]. Unlike prior approaches, we explore a temporal information infusion method to ensure spatio-temporal representation modeling in a broader range.
Token aggregation. Conventional token aggregation methods, prioritizing efficiency enhancement, categorize into hard and soft clusterings. In hard clustering methods like K-Means [21], K-Medoids [15], and DPC-KNN [12], clusters are distinct without overlap, and every data point is assigned solely to one cluster. Bipartite matching [14] divides the data points into two groups and merges a specific number of edges with the highest scores. Soft clustering allows each data point to belong to multiple clusters with varying likelihood. Fuzzy C-Means [7] extends K-Means by enabling non-exclusive partitions and prototypes. Recent token aggregation methods for efficient ViTs [19, 1, 25, 35, 29] leverage aggregation algorithms to reduce the number of tokens for efficiency. While previous approaches focus on efficiency, we employ token aggregation approaches to link relevant tokens across the videos for contextual information.
Prompt learning. Many works have proposed to transfer VLMs to downstream tasks by optimizing a discrete set of prompt vectors [11, 38, 37, 17, 26, 32, 13]. In video recognition, [13] has introduced text prompt tuning, while ViFi-CLIP [26] and Vita-CLIP [32] perform prompting in both vision and text branches. However, these prompt vectors are individually optimized and not shared across the modalities. In image recognition, Co-CoOp [37] performs an instance-conditional prompt tuning by explicitly conditioning the text prompts on [CLS] tokens from image instances. MaPLe [17] learns multi-modal prompting by sharing layerwise context prompts for both branches. Unlikely, we generate video-conditional prompts by utilizing contextualized tokens as vision inputs and injecting summarized video information into text prompt vectors.
@@ 3.1 Preliminary
3.1.1 Video encoding.
Given a video $V\in\mathbb{R}^{T\times H\times W\times 3}$ of spatial resolution $H\times W$ with $T$ sampled frames, following the Vision Transformer (ViT) architecture [6], we first divide each frame into $P\times P$ non-overlapping patches and flattening them as a set of vectors $\{\mathbf{x}_{t,i}\in\mathbb{R}^{3P^{2}}\}_{i=1}^{N}$, where $t$ is the frame index, $i$ is the patch index, and $N=HW/P^{2}$ is the number of patches. A sequence of tokens in the $t^{\text{th}}$ frame, denoted by $\mathbf{z}_{t}^{0}$, is fed into the CLIP vision encoder after combining with the spatial positional encoding $\mathbf{e}_{\text{pos}}$ as
@@
where $\mathbf{x}_{\text{cls}}\in\mathbb{R}^{d_{v}}$ is a learnable classification embedding, [CLS] token, and $\mathbf{W}_{\text{emb}}\in\mathbb{R}^{3P^{2}\times d_{v}}$ is a linear projection matrix. The CLIP vision encoder, $f_{\theta_{v}}(\cdot)$, sequentially encodes the frame-level representation $\mathbf{z}_{t}^{l}$ at each layer $l\in\{1,\ldots,L_{v}\}$ as follows:
$\mathbf{z}_{t}^{l}=f_{\theta_{v}}^{l}(\mathbf{z}_{t}^{l-1}),$ (2)
where $f_{\theta_{v}}^{l}(\cdot)$ is the $l^{\text{th}}$ layer of the CLIP vision encoder. Then, we project the [CLS] token of the $t^{\text{th}}$ frame, $\mathbf{z}_{t,0}^{L_{v}}$, onto a common vision-language latent space:
$\mathbf{v}_{t}=\mathbf{z}_{t,0}^{L_{v}}\mathbf{W}_{\text{vis}},$ (3)
where $\mathbf{W}_{\text{vis}}\in\mathbb{R}^{d_{v}\times d_{vl}}$ is a projection matrix. Finally, the video representation $\mathbf{v}$ is obtained by averaging the per-frame representations $\mathbf{v}_{t}$ as
$\mathbf{v}=\text{AvgPool}([\mathbf{v}_{1},...,\mathbf{v}_{T}]).$ (4)
3.1.2 Text encoding.
Given a text description $C$ for a video, the input of the text encoder, $\textbf{c}^{0}$, is obtained by tokenizing words in the description and computing their word embedding vectors. In addition to the embeddings from category names, one can augment a sequence of prompt embeddings $\textbf{p}^{0}$, which are obtained from either hand-crafted templates such as “a photo of a” or learnable prompt vectors. The CLIP text encoder, $f_{\theta_{c}}(\cdot)$, sequentially processes a sequence of text embeddings including prompt embeddings, which is denoted by $[\textbf{p}^{0},\textbf{c}^{0}]$, and computes an intermediate representation at each layer $l\in\{1,\ldots,L_{c}\}$ as follows:
@@
3.1.3 Video-text alignment.
Finally, the similarity between video and text embeddings are formulated as follows:
$\mathrm{sim}(\mathbf{v},\mathbf{c})=\frac{\langle\mathbf{v},{\mathbf{c}\rangle}}{\lVert\mathbf{v}\rVert\lVert\mathbf{c}\rVert}.$ (6)
During training, the goal is to maximize the similarity if $V$ and $C$ are matched but minimize it otherwise. For inference, the category with the highest similarity is chosen as the prediction output.
@@ 3.2 Temporal Contextualization (TC)
To alleviate the chronic drawbacks of existing CLIP-based video representations and leverage the spatio-temporal context in an entire video during the encoding process, we propose Temporal Contextualization (TC). TC consists of three steps, which are 1) informative token selection in each frame, 2) context summarization across spatio-temporal dimensions, and 3) temporal context infusion to all tokens in the subsequent layer. Fig. 4 illustrates an overview of TC. We discuss the three steps of our TC pipeline in the following.
Figure 4: Overview of Temporal Contextualization (TC). (a) Starting with the original patch tokens of the entire video, we first select informative seed tokens from each frame and perform temporal context summarization to aggregate relevant tokens to obtain context tokens. (b) The encoded context tokens are subsequently utilized as key-value pairs for self-attention in the next layer. (c) We employ attention score with [CLS] tokens as the criterion for seed token selection.
3.2.1 Informative token selection.
Due to a large number of redundant patches, the use of all patches for encoding a video may be suboptimal for extracting desired temporal information. Our approach focuses on selecting the informative tokens before performing the token aggregation. To this end, the attention scores obtained from self-attention operations in each frame are used as criteria to select seed tokens. Specifically, given patch tokens $\{\mathbf{z}_{t,i}\}_{i=1}^{N}$ in the $t^{\text{th}}$ frame, a set of attention scores $\{\mathbf{a}(\mathbf{z}_{t,i})\}_{i=1}^{N}$ is driven from the attention of the query of the [CLS] token to the keys of patch tokens, which is given by
@@
where both the query $\mathbf{q}_{\text{cls}}=\mathbf{z}_{\text{t,0}}\mathbf{W}_{q}\in\mathbb{R}^{d}$ and keys $\mathbf{K}_{\mathbf{z}_{t}}=\mathbf{z}_{t}\mathbf{W}_{k}\in\mathbb{R}^{(N+1)\times d}$ are given by linear projections of input $\mathbf{z}_{t}\in\mathbb{R}^{(N+1)\times d}$. In practice, our model yields multiple [CLS] attention scores through multi-head self-attention (MHSA) operations and compute the average of the attention scores from all heads, i.e., $\bar{\mathbf{a}}_{t,i}=\sum_{h=1}^{H}\mathbf{a}_{t,i}^{h}/H$, where $\mathbf{a}_{t,i}^{h}=\mathbf{a}^{h}(\mathbf{z}_{t,i})$ is the attention score for the $i^{\text{th}}$ patch $\mathbf{z}_{t,i}$ in the $t^{\text{th}}$ frame and $H$ is a number of heads. Finally, we formulate a set of seed tokens for the $t^{\text{th}}$ frame, $\mathcal{S}_{t}$, by selecting the indices for $n_{s}$ elements with highest attention score $\bar{\mathbf{a}}_{t,i}$, where the number of seed tokens in each frame $n_{s}$ is controlled by the ratio of $\alpha=n_{s}/N$.
3.2.2 Temporal context summarization.
We now interconnect the seed tokens derived from individual frames based on their relevance and identify a collection of context tokens. Specifically, we collect all the seed tokens from all frames as $\mathcal{S}=\{(t,i)|i\in\mathcal{S}_{t},t=1,\ldots,T\}$ and perform their spatio-temporal summarization by clustering and merging all the seed tokens in $\mathcal{S}$ as follows:
$\hat{\mathbf{s}}=\phi\big{(}\{\hat{\mathbf{z}}_{t,i}\}_{(t,i)\in\mathcal{S}}\big{)}.$ (8)
where $\hat{\mathbf{s}}\in\mathbb{R}^{k\times D}$ denotes a collection of the summarized tokens, which we call context tokens, $\phi$ is a token aggregation function, and $\hat{\textbf{z}}_{t,i}$ indicates an interim token encoded from $\textbf{z}_{t,i}$ via the self-attention operation. In practical, we construct a bipartite graph whose vertices and edges represent the original seed tokens and their pairwise similarities. Then, we iteratively perform bipartite matching [14] that splits the vertex set and merges $r$ pairs of vertices based on their similarities by simple averaging of vertex features until there remain $k$ number of tokens. Subsequently, the context tokens $\hat{\mathbf{s}}$ are fed into a feed-forward network (FFN) as done in the original encoding procedure and are forwarded to the next layer.
3.2.3 Temporal context infusion.
Finally, we infuse the summarized context to all patch tokens by modifying the self-attention function. The keys and values of self-attention in every frame are expanded to include context tokens as follows:
@@
where $\mathbf{K}_{\mathbf{s}}=\mathbf{s}\mathbf{W}_{k}$ and $\mathbf{V}_{\mathbf{s}}=\mathbf{s}\mathbf{W}_{v}$ are linear projections of the context tokens $\mathbf{s}\in\mathbb{R}^{k\times d}$. Here, $\mathbf{B}\in\mathbb{R}^{(N+1)\times(N+k+1)}$ is a bias matrix that distinguishes between frame-level local information and video-level global information in the expanded key matrix as
@@
where $\text{MHSA}_{\text{TC}}(\cdot,\cdot)$ denotes the TC-based MHSA operation including Eq. (9) and $\text{LN}(\cdot)$ stands for the layer normalization function.

### v2

@@ Abstract
We propose a novel framework for video understanding, called Temporally Contextualized CLIP (TC-CLIP), which leverages essential temporal information through global interactions in a spatio-temporal domain within a video. To be specific, we introduce Temporal Contextualization (TC), a layer-wise temporal information infusion mechanism for videos, which 1) extracts core information from each frame, 2) connects relevant information across frames for the summarization into context tokens, and 3) leverages the context tokens for feature encoding. Furthermore, the Video-conditional Prompting (VP) module processes context tokens to generate informative prompts in the text modality. Extensive experiments in zero-shot, few-shot, base-to-novel, and fully-supervised action recognition validate the effectiveness of our model. Ablation studies for TC and VP support our design choices. Our project page with the source code is available at https://github.com/naver-ai/tc-clip.
Keywords:
Video Action Recognition Vision-Language Model
@@ 1 Introduction
Figure 1: Comparison of attention maps between various temporal modeling approaches. Both (a) and (b) fail to recognize actions in the latter frames, whereas (c) exhibits weak discriminability due to sparse attention on the background. In contrast, (d) our method successfully focuses on informative regions across all frames, leading to the accurate action recognition result.
@@
Figure 2: Temporal information learning methods. Prior works consider temporal cues during the encoding process via (a) cross-frame attention [30, 43] with [CLS] token interactions or (b) temporal window expansion [44] by adding adjacent frame tokens to key-value pairs. However, the former lacks patch-level interactions, while the latter limits the range of temporal interactions. (c) Joint space-time attention allows full interactions across all tokens, but it is costly and suboptimal in practice (see Fig. 3.) (d) Unlike prior approaches, our method aggregates pivotal tokens from a broader range yet efficiently for enhanced temporal integration into key-value pairs.
@@
Figure 3: Pitfall of joint space-time attention. (a) Extending CLIP’s temporal sequence length degrades attention quality, presumably because it was not trained on such long sequences. (b) We compare the action recognition performance in the few-shot setting on diverse datasets. All existing methods fall behind our method.
Pretrained large-scale Vision-Language Models (VLMs) have shown remarkable generalization capability in video understanding and have emerged as promising tools even for zero-shot or open-vocabulary recognition tasks [33, 12, 50]. However, pretraining task-specific models using video-text pairs pose significant challenges, primarily due to substantial computational costs and excessive expense for annotated video-text data [42, 45]. Consequently, recent studies in video understanding [41, 30, 15, 43, 35, 44, 5, 11] have shifted their focus toward employing image-based VLMs such as CLIP [33] with fine-tuning for aligning video representations with text embeddings derived from category names.
Despite the success of CLIP in video recognition, existing approaches fail to model temporal information in the video feature learning process, as shown in Fig. 1(a)-(b). This limitation stems from the restrictive token interactions in the temporal axis. For instance, cross-frame attention approaches [30, 43], shown in Fig. 2(a), attempt to gather temporal information only through class tokens. Although VCLIP [44] incorporates patch-level details by bringing keys and values from neighborhood frames in its self-attention operation as in Fig. 2(b), its temporal scope is too narrow. Furthermore, ViFi-CLIP [35] simply averages frame-wise representations with no inter-frame information exchanges. Such naïve approaches tend to bias the models towards static information in their representation learning (e.g., objects and backgrounds) and hamper learning temporal dynamics (e.g., motion and temporal variations). To ensure the global interactions of patch tokens in a spatio-temporal domain, one possible option is to consider every patch token from all frames as a reference during the encoding process as illustrated in Fig. 2(c).
Unfortunately, such a straightforward extension for temporally global interactions in VLMs pretrained with short image-text pairs witnesses extrapolation challenges [32, 4]; we have observed that a naïve extension of sequence length along the temporal axis degrades its discriminability substantially, as shown in Fig. 3(a). The joint space-time attention model spreads attention over many patches and fails to focus on informative tokens to recognize actions, resulting in suboptimal performance compared to the frame-wise attention baseline. Moreover, this approach suffers from heavy computational overhead due to numerous redundant and similar tokens, which often correspond to background regions.
This paper presents Temporally Contextualized CLIP (TC-CLIP), a novel paradigm for extending CLIP to videos by encoding holistic video information through advanced temporal analysis. Specifically, our Temporal Contextualization (TC) pipeline summarizes global action cues into a small set of tokens, called context tokens, for reference during the encoding process. These context tokens act as additional key-value pairs for attention operations, presumably serving as temporal bridges that convey the video-level context. Our preliminary study, shown in Fig. 3(b), implies that existing methods illustrated in Fig. 2 offer minimal gains over the frame-wise attention, highlighting the need for enhanced token interactions.
@@
In addition, a Video-conditional Prompting (VP) module generates instance-level textual prompts based on context tokens from the vision encoder. The VP module comprises cross-attention operations that adopt learnable text prompts as queries and context tokens as keys and values to inject video instance representations into video-conditional textual prompts. This strategy compensates for the lack of textual semantics in action recognition datasets, where textual descriptions are limited to action class names (e.g., skateboarding, skydiving) without detailed narratives.
To verify the effectiveness and robustness of TC-CLIP, we perform extensive evaluations across diverse video recognition benchmarks. Quantitative comparisons in zero-shot, few-shot, base-to-novel, and fully-supervised settings show that the proposed approach outperforms the state-of-the-art methods by significant margins. We also provide an in-depth analysis of our design choices and the impact of each component in our model.
@@ 2.1 Preliminary
Video encoding. Suppose that there exists a video $V\in\mathbb{R}^{T\times H\times W\times 3}$ of a spatial resolution $H\times W$ with $T$ sampled frames. Following the Vision Transformer (ViT) architecture [8], we first divide each frame into $P\times P$ non-overlapping patches and flatten them as a set of vectors $\{\mathbf{x}_{t,i}\in\mathbb{R}^{3P^{2}}\}_{i=1}^{N}$, where $t$ is the frame index, $i$ is the patch index, and $N=HW/P^{2}$ is the number of patches. Then, we derive a frame-level token sequence, $\mathbf{z}_{t}^{0}$ as follows:
@@
where $\mathbf{x}_{\text{cls}}\in\mathbb{R}^{d_{v}}$ is a learnable classification embedding named [CLS] token, $\mathbf{W}_{\text{emb}}\in\mathbb{R}^{3P^{2}\times d_{v}}$ is a linear projection matrix, and $\mathbf{e}_{\text{pos}}$ is the spatial positional embedding. The CLIP vision encoder, $f_{\theta_{v}}(\cdot)$, sequentially encodes $\mathbf{z}_{t}^{l}$ at each layer $l\in\{1,\ldots,L_{v}\}$, which is given by
$\mathbf{z}_{t}^{l}=f_{\theta_{v}}^{l}(\mathbf{z}_{t}^{l-1}).$ (2)
We project the [CLS] token of the $t^{\text{th}}$ frame, denoted by $\mathbf{z}_{t,0}^{L_{v}}$, onto a common vision-language latent space using a matrix $\mathbf{W}_{\text{vis}}\in\mathbb{R}^{d_{v}\times d_{vl}}$, i.e., $\mathbf{v}_{t}=\mathbf{z}_{t,0}^{L_{v}}\mathbf{W}_{\text{vis}}$. Finally, the video representation $\mathbf{v}$ is obtained by averaging the per-frame representations $\mathbf{v}_{t}$ as $\mathbf{v}=\text{AvgPool}([\mathbf{v}_{1},...,\mathbf{v}_{T}])$.
Text encoding. Given a text description $C$—category name in our problem—for a video, the input of the text encoder, $\textbf{c}^{0}$, is obtained by tokenizing words in the description and computing their word embedding vectors. In addition to the embeddings from category names, one can augment a sequence of prompt embeddings $\textbf{p}^{0}$, which are obtained from either hand-crafted templates such as “a photo of a” or learnable prompt vectors. The CLIP text encoder, $f_{\theta_{c}}(\cdot)$, sequentially processes a sequence of text embeddings including prompt embeddings, denoted by $[\textbf{p}^{0},\textbf{c}^{0}]$, and computes an intermediate representation at each layer $l\in\{1,\ldots,L_{c}\}$ as follows:
@@
Video-text alignment. The similarity between video and text embeddings are formulated as $\mathrm{sim}(\mathbf{v},\mathbf{c})=\frac{\langle\mathbf{v},{\mathbf{c}\rangle}}{\lVert\mathbf{v}\rVert\lVert\mathbf{c}\rVert}$. During training, the goal is to maximize the similarity if $V$ and $C$ are matched and minimize otherwise. For inference, the category with the highest similarity is chosen as the prediction.
2.2 Motivation
Table 1: Motivation: What is the proper format for reference tokens? We compare 16-shot training results using various types of reference tokens during the frame-level representation encoding process. Using context tokens consistently improves the baseline model regardless of the choice of the token aggregation function.
Type of reference tokens HMDB-51 UCF-101 SSv2
Baseline (No reference tokens) 67.1 93.3 12.0
[CLS] tokens from all frames [30, 43] 67.2 (+0.1) 93.2 ($-0.1$) 12.3 (+0.3)
Patch tokens from adjacent frames [44] 67.8 (+0.7) 93.2 ($-0.1$) 12.8 (+0.8)
Patch tokens from all frames 63.3 ($-3.8$) 91.9 ($-1.4$) 12.0 (+0.0)
Context tokens — K-means [26] 68.0 (+0.9) 93.3 (+0.0) 13.1 (+1.1)
Context tokens — DPC-KNN [14] 67.9 (+0.8) 94.0 (+0.7) 14.3 (+2.3)
Context tokens — Bipartite soft matching [16, 1] 68.0 (+0.9) 93.8 (+0.5) 14.3 (+2.3)
Context tokens — Saliency-aware bipartite matching [6] 67.3 (+0.2) 93.7 (+0.4) 13.6 (+1.6)
Despite the successful generalization of CLIP for action recognition, its visual feature encoding process in Eq. (2) constrains the model’s ability to capture comprehensive spatio-temporal dynamics because it only considers intra-frame token relationships. This limitation has led previous works to additionally incorporate reference tokens, denoted by $\mathbf{s}$, to encode the $t^{\text{th}}$ frame tokens $\mathbf{z}_{t}$ as
$\mathbf{z}_{t}^{l}=f_{\theta_{v}}^{l}(\mathbf{z}_{t}^{l-1},\mathbf{s}^{l-1}).$ (4)
However, their reference token designs are still limited due to insufficient spatio-temporal interaction range. For instance, cross-frame attention (Fig. 2(a)) [30, 43] utilizes learnable global embedding vectors, e.g., [CLS] tokens, from all frames to define the reference token as $\mathbf{s}=[\mathbf{z}_{1,0},...,\mathbf{z}_{T,0}]$, and temporal window expansion (Fig. 2(b)) [44], on the other hand, integrates neighboring frame patch tokens for $\mathbf{s}=[\mathbf{z}_{t-1},\mathbf{z}_{t+1}]$. Note that the former lacks patch-level details whereas the latter captures temporal information only within a local range. Although incorporating all patch tokens from a whole video as a reference (Fig. 2(c)), $\mathbf{s}=[\mathbf{z}_{1},...,\mathbf{z}_{T}]$, might be conceptually reasonable, it is not practical due to the excessive number of tokens. Furthermore, this approach conflicts with the properties of CLIP pretrained on short image-text pairs and significantly degrades attention quality.
To this end, we compute a reference, $\mathbf{s}=\phi([\mathbf{z}_{1},...,\mathbf{z}_{T}])$, using a small set of context tokens that summarize a whole input video, where $\phi(\cdot)$ is a token aggregation function. This approach delivers the essential temporal information for the feature encoding while maintaining CLIP’s effective sequence length. Our preliminary study, presented in Table 1, shows that using the proposed context tokens as a reference consistently outperforms the frame-wise attention baseline, whereas all other types of reference tokens only yield marginal gains or even suffer from performance degradation.
@@ 2.3 Temporal Contextualization (TC)
Figure 4: Overview of Temporal Contextualization (TC). We first collect informative tokens from each frame and then aggregate relevant seed tokens to obtain context tokens. They are used as key-value pairs for the self-attention in the next layer.
Based on our motivation, we propose Temporal Contextualization (TC), which consists of three steps: 1) informative token selection in each frame, 2) context summarization across spatio-temporal dimensions, and 3) context infusion to all tokens in the subsequent layer. Fig. 4 illustrates an overview of TC.
Informative token selection. Due to the many redundant tokens in a video, using all patches may be suboptimal for extracting desired temporal information. We only select the informative seed tokens based on each frame’s attention scores obtained from self-attention operations. Specifically, given patch tokens $\{\mathbf{z}_{t,i}\}_{i=1}^{N}$ in the $t^{\text{th}}$ frame, a set of attention scores $\{\mathbf{a}(\mathbf{z}_{t,i})\}_{i=1}^{N}$ is driven from the attentiveness of the [CLS] token with respect to other tokens, which is given by
@@
where both the query $\mathbf{q}_{\text{cls}}=\mathbf{z}_{\text{t,0}}\mathbf{W}_{q}\in\mathbb{R}^{d}$ and keys $\mathbf{K}_{\mathbf{z}_{t}}=\mathbf{z}_{t}\mathbf{W}_{k}\in\mathbb{R}^{(N+1)\times d}$ are given by linear projections of input $\mathbf{z}_{t}\in\mathbb{R}^{(N+1)\times d}$. In practice, our model yields multiple [CLS] attention scores from multi-head self-attention (MHSA) operations and computes the average of the attention scores from all heads, i.e., $\bar{\mathbf{a}}_{t,i}=\sum_{h=1}^{H}\mathbf{a}_{t,i}^{h}/H$, where $\mathbf{a}_{t,i}^{h}=\mathbf{a}^{h}(\mathbf{z}_{t,i})$ is the attention score for the $i^{\text{th}}$ patch $\mathbf{z}_{t,i}$ in the $t^{\text{th}}$ frame and $H$ is the number of heads. Finally, we identify a set of seed token indices for the $t^{\text{th}}$ frame, $\mathcal{S}_{t}$, by selecting $n_{s}$ elements with the highest attention scores, where $n_{s}$ is controlled by a hyperparameter $\alpha=n_{s}/N$.
Temporal context summarization. We describe how to connect the seed tokens derived from individual frames based on their relevance and identify a collection of context tokens. We first collect the seed tokens from all frames as $\{\hat{\mathbf{z}}_{t,i}\}_{(t,i)\in\mathcal{S}}$, where $\mathcal{S}=\{(t,i)|i\in\mathcal{S}_{t},t=1,\ldots,T\}$ is a set of seed token indices from all frames and $\hat{\textbf{z}}_{t,i}$ indicates an interim token encoded from $\textbf{z}_{t,i}$ via the self-attention operation. Then we perform their spatio-temporal summarization by clustering and merging all the seed tokens as
$\hat{\mathbf{s}}=\phi\big{(}\{\hat{\mathbf{z}}_{t,i}\}_{(t,i)\in\mathcal{S}}\big{)},$ (6)
where $\hat{\mathbf{s}}\in\mathbb{R}^{k\times D}$ denotes a collection of the summarized tokens, which we call context tokens, and $\phi$ is a token aggregation function. While diverse token aggregation techniques are valid for TC (See Table 8), we adopt bipartite soft matching [16, 1] by default. Subsequently, the context tokens $\hat{\mathbf{s}}$ are fed into a feed-forward network (FFN).
Temporal context infusion. Finally, we infuse the summarized context to all patch tokens by modifying the self-attention function. The keys and values of self-attention in every frame are expanded to include context tokens as follows:
@@
where $\mathbf{K}_{\mathbf{s}}=\mathbf{s}\mathbf{W}_{k}$ and $\mathbf{V}_{\mathbf{s}}=\mathbf{s}\mathbf{W}_{v}$ are linear projections of the context tokens $\mathbf{s}\in\mathbb{R}^{k\times d}$. Here, $\mathbf{B}\in\mathbb{R}^{(N+1)\times(N+k+1)}$ is a bias matrix that distinguishes between frame-level local information and video-level global information in the expanded key matrix as follows:
@@
where $\text{MHSA}_{\text{TC}}(\cdot,\cdot)$ denotes the MHSA operation based on Eq. (7) and $\text{LN}(\cdot)$ stands for the layer normalization function.